Calcium hydride, CaH2, reacts with water to form hydrogen gas:

CaH2 (s) + 2H2O (l) → Ca(OH)2 (aq) + 2H2 (g)

The reaction is sometimes used to inflate life rafts, weather balloons, and the like, where a simple, compact means of generating H2 is desired. How many grams of CaH2 are needed to generate 64.5 L of H2 gas if the pressure of H2 is 814 torr at 32 oC?


1
Expert's answer
2022-04-25T14:42:04-0400

Solution:

Convert °C to K:

32°C + 273.15 = 305.15 K


Calculate the moles of hydrogen gas (H2) using the ideal gas equation (PV = nRT):

n(H2) = PV / RT

n(H2) = (814 torr × 64.5 L) / (62.36 L torr mol−1 K−1 × 305.15 K) = 2.759 mol

Moles of H2 = 2.759 mol


Balanced chemical equation:

CaH2(s) + 2H2O(l) → Ca(OH)2(aq) + 2H2(g)

According to stoichiometry:

1 mol of CaH2 produces 2 mol of H2

X mol of CaH2 produces 2.759 mol of H2

Thus,

Moles of CaH2 = X = (2.759 mol H2) × (1 mol CaH2 / 2 mol H2) = 1.3795 mol CaH2


The molar mass of CaH2 is 42.094 g/mol

Therefore,

Mass of CaH2 = (1.3795 mol CaH2) × (42.094 g CaH2 / 1 mol CaH2) = 58.07 g CaH2 = 58.1 g CaH2

Mass of CaH2 = 58.1 g


Answer: 58.1 grams of CaH2 are needed